The tub of a washer goes into its spin cycle, starting from rest and gaining angular speed steadily for 18 s, at which time it is turning at 11 rev/s. At this point the person doing the laundry opens the lid, and a safety switch turns off the washer. The tub smoothly slows to rest in 12 s. Through how many revolutions does the tub turn while it is in motion? Encode only the numerical value. Whole number.
Answer #1

Initial speed of the tub = 0

Final speed of the tub = 11.0 rev/s

Since the acceleration is constant, the speed profile is linear, so just use the average speed and multiply by time.

The average speed up, and average speed down, is:

(1/2)(0+11.0)rev/s = 5.5 rev/s

Revolutions speeding up:

18.0 s x 5.5 rev/s = 99.0 rev

Revolutions slowing down:

12.0 s x 5.5 rev/s = 66.0 rev

Therefore, total revolutions while in motion = 99.0 + 66.0 = 165.0 revolutions. (Answer)